Two resistances of 10 and 15 ohms each respectively are connected in parallel. The two are then connected in series with a 5-ohm resistance. What is the equivalent resistance?

1
Expert's answer
2022-01-26T17:46:25-0500

Given R1 = 10 Ohms

R2 = 15 Ohms

R3 = 5 Ohms



R1 and R2 are in parallel, therefore, R12 = R1 || R2

R12 = "\frac{1}{R12} = \frac{1}{R1} + \frac{1}{R2}"

R12 = "\frac{R1 * R2}{R1+R2}"

R12 = "\frac{10 + 15}{10*15} = 6 Ohms"


Now R12 and R3 are in series,

Therefore, Net equivalent Res. R = R12 + R3 = 6 + 5 = 11 Ohms

Therefore, Final Equivalent Res. = 11 Ohms